OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. y + 2 = 1/3(x – 5); (–4, 3) (1 point)

OpenStudy (anonymous):

***Or at least help me please (:

OpenStudy (amistre64):

need to be able to recognize the slope of the given line. that will be helpful to determine the perp slope

OpenStudy (anonymous):

Thank you

OpenStudy (amistre64):

the general line form you have given is: \[y-y_o=m(x-x_o)\] where m is the slope

OpenStudy (amistre64):

a perp line has the negative flip of the given m; can you define the "m" part for me?

OpenStudy (anonymous):

1/3? I'm really not all that good at this

OpenStudy (amistre64):

you are correct :) the negative flip of that is what we want: 1/3 flips and negates to -3

OpenStudy (amistre64):

so we have at least: \[y-y_o=-3(x-x_o)\] the x0 and yo parts are the given point (-4,3) so they just fill in can you show me the finished results?

OpenStudy (anonymous):

y - -4 = -3(x -3) Sorry if I'm way off

OpenStudy (amistre64):

you did that perfectly; and to clean it up, remember that 2 negatives is a positive: -- = + y + 4 = -3(x-3) good job

OpenStudy (amistre64):

wait ... almost perfectly

OpenStudy (amistre64):

you got the x and y inthe wrong spots, and i did too

OpenStudy (amistre64):

xo = -4, yo= 3 y -3 = -3(x -- 4) y -3 = -3(x + 4) thats better

OpenStudy (anonymous):

Oh okay thank you so much

Parth (parthkohli):

We start a letter with "yo" and end it with "xo" :-)

OpenStudy (amistre64):

if you are satisfied with the results of this question, go ahead and close it so that you can ask a new question. Otherwise people wont really know if you are asking a new one, or replying to this old one

OpenStudy (anonymous):

How do I close it? sorry

OpenStudy (anonymous):

Oh wait I figured it out. Thank you!

OpenStudy (amistre64):

over on the left side of the screen it says bump close; just hit close

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

good luck ;)